OpenStudy (anonymous):

One number is three less than four times another. If their difference is 48, what is the larger number?

OpenStudy (bahrom7893):

x = 4y-3 x-y = 48

OpenStudy (anonymous):

you jus lost me?

OpenStudy (anonymous):

Bahrom just wrote two expressions describing what you said in English. To solve it, find what x or y equals and plug that value into the other expression. I.e., x=4y-3 x-y=4y-3-y=48 4y-y-3=48 = 3y-3=48 3y=51 y=51/3 x=4(51/3)-3

OpenStudy (bahrom7893):

yes.. just finish solving..

OpenStudy (anonymous):

51/3 are we dividin?

OpenStudy (bahrom7893):

yea..